OpenStudy (anonymous):

I need help with this; Steven wants to buy some candy with his birhtday money. He can choose from Jelly Belly jelly beans that sell for 6.99 per pound and Barch's variety that sells for 3.99. He likes to have twice the amount of jelly beans as Brachs variety. If he spent a total of 53.91 , how many pounds of each type of candy did he buy?

OpenStudy (anonymous):

x=jelly belly y=Brachs there's 2 times as many JB as there is Brachs. x=2y 53.91=6.99x+3.99y Substitute 2y for x in the second equation. 53.91=6.99(2y)+3.99y Solve that. 53.91=13.98y+3.99y 53.91=17.97y 3=y Substitute 3 for y in the first equation. x=2(3) x=6 So he bought 3 Brachs and 6 Jelly Belly